### 28 Feb 2024 v6

# The shocks in Josephson transmission line revisited

Abstract
We continue our previous studies of the shocks in the lossy Josephson transmission line (JTL). The paper consists of two parts. In the first part we analyse the scattering of the ”sound’ (small amplitude small wave vector harmonic wave) on the shock wave and calculate the reflection and the transmission coefficients. In the second part we show that the kinks, which we previously studied only in the lossless JTL, exist also in the lossy JTL and study the similarities and the dissimilarities between the shocks and the kinks there. We find that the nonlinear equation describing the weak kinks and the weak shocks can be integrated (in particular cases) in terms of elementary functions. We also show that the profile of the shock in the lossy JTL demonstrates oscillatory behavior with leading peaks resembling solitary waves if the losses are weak.
I Introduction
The interest in studies of nonlinear electrical transmission lines, in particular of lossy nonlinear transmission lines, has started some time ago rosenau ; chen ; mohebbi , but it became even more pronounced recently ricketts ; houwe ; katayama ; sekulic . In ref. malomed2 , one can find a very recent and complete review of studies of nonlinear electric transmission networks.
We studied previously the shock waves in the lossy Josephson transmission line (JTL) JTL kogan1 ; kogan2 ; kogan3 and kinks and solitons in the lossless (actually, without any shunting at al) JTL kogan2 ; kogan3 . The present work had several aims. First we would like to analyse the interaction between the ”sound” (small amplitude small wave vector harmonic wave) and the shock wave. Second we would like to establish the relation between the shock waves and the kinks. And third, we would like to additionally study the weak waves, and, in particular, to look for the cases when the nonlinear equation, describing such waves in the JTL, can be integrated in terms of elementary functions.
The rest of the article is constructed as follows. In Section II we rederive the circuit equations describing the JTL in the continuum approximation. In Section III we consider scattering of the ”sound” wave by the shock wave and calculate the appropriate reflection and transmission coefficients. In Section IV we show that the kinks which represent stationary solutions in the lossless JTL, can be also observed as weakly attenuating disturbances in the lossy JTL, which support stationary shock waves. The connection between the shocks and the kinks in lossy JTL is revealed. We show that the profile of the shocks for the case of weak losses demonstrates soliton-like features. We also integrate the wave equation describing weak shocks (and kinks) in terms of elementary functions for the specific value of the losses parameter. We conclude in Section V. In the Appendix A we present a physically appealing model of the JTL, composed of superconducting grains. In the Appendix B we formulate the condition for the applicability of the continuum approximation used in the paper. Some mathematical details are relegated to Appendices C and D.
II The circuit equations: continuum approximation
The discrete model of the Josephson transmission line (JTL) is constructed from the identical Josephson junctions (JJs) capacitors and resistors, as shown on Fig. 1. (Possible physical realization of the model is presented in the Appendix A.) We take as the dynamical variables the phase differences (which we for brevity will call just phases) $\varphi_{n}$ across the JJs and the voltages $v_{n}$ of the ground capacitors. The circuit equations are
$\frac{\hbar}{2e}\frac{d\varphi_{n}}{dt}$ $=v_{n-1}-v_{n}$ (1a)
$C\frac{dv_{n}}{dt}$ $=I_{c}\sin\varphi_{n}-I_{c}\sin\varphi_{n+1}$
$+\left(\frac{1}{R_{J}}+C_{J}\frac{d}{dt}\right)\frac{\hbar}{2e}\frac{d}{dt}\left(\varphi_{n}-\varphi_{n+1}\right),$ (1b)
where $C$ is the capacitance, $I_{c}$ is the critical current of the JJ, and $C_{J}$ and $R_{J}$ are the capacitor and the ohmic resistor shunting the JJ.
Figure 1: Discrete Josephson transmission line.
In the continuum approximation we treat $n$ as the continuous variable $Z$ and approximate the finite differences in the r.h.s. of the equations by the first derivatives with respect to $Z$, after which the equations take the form
$\frac{\partial\varphi}{\partial\tau}$ $=-\frac{\partial V}{\partial Z},$ (2a)
$\frac{\partial V}{\partial\tau}$ $=-\frac{\partial}{\partial Z}\left(\sin\varphi+\frac{Z_{J}}{R_{J}}\frac{\partial\varphi}{\partial\tau}+\frac{C_{J}}{C}\frac{\partial^{2}\varphi}{\partial\tau^{2}}\right).$ (2b)
where we introduced the dimensionless time $\tau=t/\sqrt{L_{J}C}$ and the dimensionless voltage $V=v/(Z_{J}I_{c})$; $L_{J}\equiv\hbar/(2eI_{c})$ is the ”inductance” of the JJ and $Z_{J}\equiv\sqrt{L_{J}/C}$ is the ”characteristic impedance” of the JTL. The condition for the applicability of the continuum approximation is formulated explicitly in the Appendix B.
III The sound scattering by the shock wave
III.1 The sound waves and the shock waves
Because (2b) is nonlinear, the system (II) has a lot of different types of solutions. In this Section we’ll be interested in only two types of those. First type - small amplitude small wave vector harmonic waves on a homogeneous background $\varphi_{0}$. For such waves Eq. (2b) is simplified to
$\frac{\partial V}{\partial\tau}=-\cos\varphi_{0}\frac{\partial\varphi}{\partial Z}.$ (3)
We ignored the shunting terms in r.h.s. of (2a) because they contain higher order derivatives in comparison with the main term, and small wave vector means also small frequency.
The harmonic wave solutions of Eq. (II) (which, for brevity we’ll call the sound) are
$\varphi$ $=\varphi_{0}+\varphi^{(h)}e^{ikz-i\omega\tau},$ (4a)
$V$ $=V_{0}+V^{(h)}e^{ikz-i\omega\tau},$ (4b)
where
$\omega=\overline{u}\left(\varphi_{0}\right)k,\hskip 28.45274pt\text{and}\hskip 28.45274pt\overline{u}^{2}\left(\varphi_{0}\right)=\cos\varphi_{0},$ (5)
$\overline{u}\left(\varphi_{0}\right)$ being the normalized sound speed. In this paper the normalized speed $\equiv$ physical speed times $\sqrt{L_{J}C}/\Lambda$, where $\Lambda$ is the JTL period. Note that the stability of a homogeneous background $\varphi_{0}$ demands
$\cos\varphi_{0}>0.$ (6)
The second type of solutions we’ll be (mostly) interested in, is shock waves kogan1 ; kogan2 , which are (locally) the solutions satisfying the conditions
$\varphi(\tau,Z)=\varphi(\tau-Z/\overline{U}),\hskip 14.22636ptV(\tau,Z)=V(\tau-Z/\overline{U}).$ (7)
Substituting the ansatz (7) into (II) and integrating thus obtained equations with respect to $\tau$ from $\infty$- ∞ to +$\infty$+ ∞ we obtain
$\overline{U}\left(\varphi_{2}-\varphi_{1}\right)$ $=V_{2}-V_{1},$ (8a)
$\overline{U}\left(V_{2}-V_{1}\right)$ $=\sin\varphi_{2}-\sin\varphi_{1},$ (8b)
where $\varphi_{1}$ and $V_{1}$ are the phase and the voltage before the shock, $\varphi_{2}$ and $V_{2}$ after the shock, and $\overline{U}$ is the normalized shock wave speed. Equation (III.1) is actually the Rankine-Hugoniot condition. The obvious result of (III.1) is:
$\overline{U}_{\varphi_{2},\varphi_{1}}^{2}=\frac{\sin\varphi_{1}-\sin\varphi_{2}}{\varphi_{1}-\varphi_{2}}.$ (9)
Note that the shunting of the JJ doesn’t influence the shock speed kogan1 ; kogan2 but determines (as we will see explicitly in the next Section) the structure of the shock front.
In this Section we ignore the structure of the shock wave and consider it as the discontinuities of the dynamical variables. Equations (III.1) connect these discontinuities with the shock speed.
III.2 The reflection and the transmission coefficients
In this Section we’ll be interested in two problems landau . The first one: A sound wave is incident from the rear on a shock wave. Determine the sound reflection coefficient $R$. The situation is shown in Fig. 2.
Figure 2: Reflection of a sound wave from a shock wave. The horizontal axis is the coordinate $Z$, the vertical axis - instantaneous value of the Josephson phase.
The second problem: A sound wave is incident from the front on a shock wave. Determine the sound transmission coefficient $T$. The situation is shown in Fig. 3.
Figure 3: Transmission of a sound wave through a shock wave. The horizontal axis is the coordinate $Z$, the vertical axis - instantaneous value of the Josephson phase.
While formulating both problems we took into account the equation, which will be derived in Section IV
$\overline{u}_{b}^{2}<\overline{U}_{\varphi_{a},\varphi_{b}}^{2}<\overline{u}_{a}^{2}.$ (10)
where $\varphi_{b}$ and $\varphi_{a}$ are the phases before and after the shock in the absence of the sound respectively. Also,
$\overline{U}_{\varphi_{2},\varphi_{1}}$ $=\overline{U}_{\varphi_{a},\varphi_{b}}+\delta\overline{U}^{(r,t)}.$ (11)
For the first problem mentioned above we have
$\varphi_{1}$ $=\varphi_{b},$ (12a)
$V_{1}$ $=V_{b},$ (12b)
$\varphi_{2}$ $=\varphi_{a}+\varphi^{(in)}+\varphi^{(r)},$ (12c)
$V_{2}$ $=V_{a}+V^{(in)}+V^{(r)},$ (12d)
where (in) stands for the incident sound wave and (r) for the reflected sound wave. Substituting (11) - (12d) into (8a), (8b) in the first order approximation we obtain
$\delta\overline{U}\left(\varphi_{a}-\varphi_{b}\right)+\overline{U}\left(\varphi^{(in)}+\varphi^{(r)}\right)=V^{(in)}+V^{(r)},$ (13a)
$\delta\overline{U}\left(V_{a}-V_{b}\right)+\overline{U}\left(V^{(in)}+V^{(r)}\right)=\overline{u}^{2}(\varphi_{a})\left(\varphi^{(in)}+\varphi^{(r)}\right).$ (13b)
Taking into account the relations
$V^{(in)}$ $=\overline{u}\left(\varphi_{a}\right)\varphi^{(in)},$ (14a)
$V^{(r)}$ $=-\overline{u}\left(\varphi_{a}\right)\varphi^{(r)}$ (14b)
(the difference in the signs is because of the opposite directions of propagation of the two waves) and excluding $\delta\overline{U}$ we obtain
$R\equiv\frac{\varphi^{(r)}}{\varphi^{(in)}}=-\frac{\left[\overline{u}\left(\varphi_{a}\right)-\overline{U}\right]^{2}}{\left[\overline{u}\left(\varphi_{a}\right)+\overline{U}\right]^{2}}=-\frac{\overline{u}_{in}^{2}}{\overline{u}_{r}^{2}},$ (15)
where $\overline{u}_{in}=\overline{u}\left(\varphi_{a}\right)-\overline{U}$ is the speed of the incident sound wave relative to the shock wave, and $\overline{u}_{r}=\overline{u}\left(\varphi_{a}\right)+\overline{U}$ is the speed of the reflected sound wave relative to the shock wave. As one could have expected, the modulus of the sound reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $\varphi_{a}\to\varphi_{b}$, in other words, when the shock wave itself nearly becomes the sound wave.
Now let us turn to the second problem. We have
$\varphi_{1}$ $=\varphi_{b}+\varphi^{(in)},$ (16a)
$V_{1}$ $=V_{b}+V^{(in)},$ (16b)
$\varphi_{2}$ $=\varphi_{a}+\varphi^{(t)},$ (16c)
$V_{2}$ $=V_{a}+V^{(t)},$ (16d)
where (t) stands for the transmitted wave. Substituting (11), (16a) - (16d) into (8a), (8b), in the first order approximation we obtain
$\delta\overline{U}\left(\varphi_{a}-\varphi_{b}\right)+\overline{U}\left(\varphi^{(t)}-\varphi^{(in)}\right)=V^{(t)}-V^{(in)}$ (17a)
$\delta\overline{U}\left(V_{a}-V_{b}\right)+\overline{U}\left(V^{(t)}-V^{(in)}\right)$
$=\overline{u}^{2}(\varphi_{a})\varphi^{(t)}-\overline{u}^{2}(\varphi_{b})\varphi^{(in)}.$ (17b)
Taking into account the relations
$V^{(in)}$ $=-\overline{u}\left(\varphi_{b}\right)\varphi^{(in)},$ (18a)
$V^{(t)}$ $=-\overline{u}\left(\varphi_{a}\right)\varphi^{(t)}$ (18b)
and excluding $\delta\overline{U}$ we obtain
$T\equiv\frac{\varphi^{(t)}}{\varphi^{(in)}}=\frac{\left[\overline{u}\left(\varphi_{b}\right)+\overline{U}\right]^{2}}{\left[\overline{u}\left(\varphi_{a}\right)+\overline{U}\right]^{2}}=\frac{\overline{u}_{in}^{2}}{\overline{u}_{t}^{2}},$ (19)
where $\overline{u}_{in}=\overline{u}\left(\varphi_{b}\right)+\overline{U}$ is the speed of the incident sound wave relative to the shock wave, and $\overline{u}_{t}=\overline{u}\left(\varphi_{a}\right)+\overline{U}$ is the speed of the transmitted sound wave relative to the shock wave. As one could have expected, the sound transmission coefficient is less than one, and goes to one when the intensity of the shock wave decreases, that is when $\varphi_{a}\to\varphi_{b}$.
Looking back at the derivation of (15) and (19) we understand that the equations will be valid also for a generalized Josephson law for the supercurrent $I_{s}$:
$I_{s}=I_{c}f(\varphi).$ (20)
where $f$ is a (nearly) arbitrary function. The difference from the case considered above is that the sound speed in the general case is
$\overline{u}^{2}\left(\varphi_{0}\right)=f^{\prime}(\varphi_{0}),$ (21)
and the shock speed is given by the equation
$\overline{U}_{\varphi_{2},\varphi_{1}}^{2}=\frac{f(\varphi_{1})-f(\varphi_{2})}{\varphi_{1}-\varphi_{2}}.$ (22)
IV The shocks and the kinks
IV.1 The travelling waves
In this Section we would like to study the structure of the shock wave, so we return to Eq. (II). Consider a solution which for $\tau\in(-\infty,+\infty)$ stays in the finite region of the phase space. The limit cycles are excluded for our problem, and strange attractors are excluded in a 2d phase space in general strogatz . Hence the trajectory begins in a fixed point and ends in a fixed point
$\lim_{\tau\to-\infty}\varphi$ $=\varphi_{1},\hskip 28.45274pt\lim_{\tau\to+\infty}\varphi=\varphi_{2}.$ (23)
In this Section we will use the ansatz (7) globally, i.e. as describing the travelling wave, $\overline{U}$ being its speed. Using the ansatz we can write down Eqs. (II) as
$\overline{U}\frac{d\varphi}{d\tau}$ $=\frac{dV}{d\tau},$ (24a)
$\overline{U}\frac{dV}{d\tau}$ $=\frac{d}{d\tau}\left(\sin\varphi+\frac{Z_{J}}{R_{J}}\frac{d\varphi}{d\tau}+\frac{C_{J}}{C}\frac{d^{2}\varphi}{d\tau^{2}}\right).$ (24b)
Note that from Eq. (24a) follows that in the travelling wave the voltage $V$ is connected to the Josephson phase $\varphi$ in a very simple way
$V=\overline{U}\varphi$ (25)
(of course, an arbitrary constant can be added to the r.h.s. of (25)).
Excluding $V$ from (IV.1) we obtain
$\overline{U}^{2}\frac{d\varphi}{d\tau}=\frac{d}{d\tau}\left(\sin\varphi+\frac{Z_{J}}{R_{J}}\frac{d\varphi}{d\tau}+\frac{C_{J}}{C}\frac{d^{2}\varphi}{d\tau^{2}}\right).$ (26)
Integrating (26) we get
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}+\sin\varphi=\overline{U}^{2}\varphi+F,$ (27)
where $\tilde{\tau}\equiv\tau\sqrt{C/C_{J}}=t/\sqrt{L_{J}C_{J}}$, $\gamma\equiv\sqrt{L_{J}/C_{J}}\big{/}R_{J}$ is the damping coefficient, and $F$ is the constant of integration. Equation (27) reminds the equation
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}+\sin\varphi=I/I_{c},$ (28)
describing current biased JJ within the RCSJ model tinkham .
IV.2 The kinks vs. the shocks
Taking into account the boundary conditions (23) and the equation for the shock speed (9), we obtain that $F$ in Eq. (27) is
$F=\frac{\varphi_{1}\sin\varphi_{2}-\varphi_{2}\sin\varphi_{1}}{\varphi_{1}-\varphi_{2}}.$ (29)
Hence Eq. (27) can be written down as
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}=-\frac{d\Pi(\varphi)}{d\varphi},$ (30)
where
$\Pi(\varphi)=\frac{\left(\varphi-\varphi_{1}\right)^{2}\sin\varphi_{2}-\left(\varphi-\varphi_{2}\right)^{2}\sin\varphi_{1}}{2(\varphi_{1}-\varphi_{2})}-\cos\varphi.$ (31)
Equation (30) is Newton equation, describing motion with friction of the fictitious particle in the potential well $\Pi(\varphi)$. The motion starts at $\tilde{\tau}=-\infty$ at $\varphi=\varphi_{1}$ and ends at $\tilde{\tau}=+\infty$ at $\varphi=\varphi_{2}$, and
$\Pi(\varphi_{2})<\Pi(\varphi_{1}).$ (32)
Because of the invariance of the system when all phases are shifted by $2\pi$, it is enough to consider $\varphi_{1}\in(-\pi,\pi)$. Because of the condition (6), we should consider only $\varphi_{1}\in(-\pi/2,\pi/2)$, and because the physics is obviously symmetric with respect to simultaneous inversion of all phases $\varphi\to-\varphi$, in the following we consider explicitly (everywhere apart from Fig. 7) only $\varphi_{1}\in(0,\pi/2)$.
We consider in this paper only the case $\varphi_{2}\in(-\pi/2,\pi/2)$. Because the sine function monotonically increases between $-\pi/2$ and $\pi/2$, the r.h.s. of (9) is always positive in this case. From (32) follows $-\varphi_{1}<\varphi_{2}<\varphi_{1}$.
If $\varphi_{2}$ is positive, it is inevitably the point of a minimum of the potential. In fact, the stationary points of the potential are given by the equation
$\sin\varphi=\overline{U}^{2}\varphi+F.$ (33)
Because $\sin\varphi$ is concave downward for $20<\varphi<\pi/$ 2, the straight line, crossing the sine curve at the points $\pi/2>\varphi_{1},\varphi_{2}>0$ can’t cross the curve in between. Hence there are no stationary points between $\varphi_{1}$ and $\varphi_{2}$. The potential $\Pi(\varphi)$ for positive $\varphi_{2}$ is illustrated in Fig. 4.
[…]
On the other hand, for $\varphi_{2}<0$ the potential $\Pi(\varphi)$ can have either a minimum or a maximum at $\varphi_{2}$, as it is illustrated in Fig. 5.
[…]
Looking at Fig. 5 (left) we realize that for the solution with $\varphi_{1}$ and $\varphi_{2}$ having opposite signs to exist, the effective friction coefficient $\gamma$ should be large enough to prevent escape of the particle above the potential barrier to the left of $\varphi_{2}$. (There is no such restriction for the shock wave with $\varphi_{1}$ and $\varphi_{2}$ having the same sign, because in this case the left potential barrier is higher than the right one, as it is illustrated in Fig. 4.)
The minimum of the potential at $\varphi_{2}$ situation corresponds to the shock wave and was discussed at length in our previous publications kogan1 ; kogan2 . In Appendix C we explain how Eq. (27) in this case can be integrated numerically. The result of such integration is presented in Fig. 6.
[…]
We considered previously the potential maximum at $\varphi_{2}$ case only in the absence of shunting kogan2 . We called such travelling waves the kinks. Now we understand that similar kinks exist also in the lossy JTL (for $-\varphi_{1}<\varphi_{2}<\varphi_{1}/2$). Looking at Fig. 5 (right), presenting the potential for the kink, we realize, that since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $\varphi_{1}$, $\varphi_{2}$ and $\gamma$ should satisfy definite relation. For weak kinks such relation will be obtained in section IV.5. Thus in the absence of losses ($\gamma=0$), only the kinks with $\varphi_{2}=-\varphi_{1}$, are possible kogan2 .
We can find the boundary between the two cases considered above (when $\varphi_{2}$ is an inflexion point of the potential) by equating the second derivative of the potential at the point $\varphi_{2}$ to zero
$\overline{U}^{2}-\overline{u}^{2}\left(\varphi_{2}\right)=0.$ (34)
The approximate solution of (34) is $\varphi_{2}=-\varphi_{1}/2$.
Everywhere above we considered the travelling wave going to the right, but, of course, by interchanging $\varphi_{1}$ and $\varphi_{2}$ (also in the inequality (32)) we obtain the wave going to the left. So the conditions for the shocks and for the kinks in the whole phase plane of the boundary conditions $(\varphi_{1},\varphi_{2})$ are shown in Fig. 7. Two additional straight lines on this Figure $\varphi_{2}=-\varphi_{1}$ and $\varphi_{2}=\varphi_{1}$ present the kinks and the solitons respectively, which can exist in the bare-bones (unshunted) JTL kogan2 and propagate in both directions.
Figure 7: The phase plane of the boundary conditions $(\varphi_{1},\varphi_{2})$. Blue regions correspond to the shock wave moving to the right, green regions - to the left. Yellow regions correspond to the kink moving to the right, red regions - to the left. The thick black line $\varphi_{2}=-\varphi_{1}$ corresponds to the kink, the thick black line $\varphi_{2}=\varphi_{1}$ to the soliton which can exist only in the bare-bones JTL and propagate in both directions.
Following the venerable tradition to tell the same story twice, we will now present an approach to integration of the equation (26), alternative to that we used above. Introducing the voltage at the Josephson junction
$E=\frac{d\varphi}{d\tilde{\tau}}$ (35)
as the new dependent variable, considering $\varphi$ as the independent variable and dividing both parts by the common multiplier $E$, we reduce the order of Eq. (26) and write it down as
$\frac{1}{2}\frac{d^{2}E^{2}}{d\varphi^{2}}+\gamma\frac{dE}{d\varphi}=\overline{U}^{2}-\overline{u}^{2}(\varphi).$ (36)
This equation will be used in Section IV.5 to study weak kinks(shocks).
Note that in the framework of the quasi-continuum approximation E is (up to a numerical multiplier) just the voltage on the JJ. Thus Eq. (36) gives the voltage on the JJ as the function of the Josephson phase.
IV.3 The shocks speed vs. the kinks speed
For the shock $\varphi_{1}$ is the point of a maximum of $\Pi(\varphi)$ and $\varphi_{2}$ is the point of a minimum. Hence the second derivative of the potential with respect to $\varphi$ is negative at $\varphi_{1}$ and positive at $\varphi_{2}$. Taking into account Eq. (9), we obtain
$\overline{u}^{2}(\varphi_{2})>\overline{U}^{2}_{\varphi_{2},\varphi_{1}}>\overline{u}^{2}(\varphi_{1}).$ (37)
The inequalities (37) reflect the well-known in the nonlinear waves theory fact: the shock speed is smaller than the sound speed in the region behind the shock but larger than the sound speed in the region before the shock whitham .
From the inequalities (37) follows that a shock can not split into two shocks moving in the same direction. Actually we can make even stronger statement: two shocks moving in the same direction will merge. In fact, let there is the first shock $\varphi_{0}\leftarrow\varphi_{1}$ and the second shock $\varphi_{2}\leftarrow\varphi_{0}$ behind it. Because of inequalities (37) the speed of the second shock is larger, and the speed of the first shock is smaller than $\overline{u}(\varphi_{0})$. The statement is proved. Note that due to a one-dimensional nature of our problem we don’t have to consider the corrugation instability of the shock wave dyakov1 ; lubchich ; semenko ; robinet ; bates .
On the other hand, two shocks going in the opposite directions may well coexist. From that we understand that any given boundary conditions $\varphi_{1},\varphi_{2}$ are compatible both with a single shock (kink) $\varphi_{2}\leftarrow\varphi_{1}$ and with two shocks $\varphi_{0}\leftarrow\varphi_{1}$ and $\varphi_{2}\rightarrow\varphi_{0}$ (with any $\varphi_{0}$, satisfying the conditions $|\varphi_{3}|<|\varphi_{1}|,|\varphi_{2}|$) going in the opposite directions. The solution with $\varphi_{0}=0$ looks especially appealing. This dichotomy will (hopefully) be analyzed in the next paper.
For the kink both $\varphi_{1}$ and $\varphi_{2}$ are the points of minima. Hence the second derivative of the potential with respect to $\varphi$ is positive at both points. Thus
$\overline{U}^{2}_{\varphi_{2},\varphi_{1}}>\overline{u}^{2}(\varphi_{2})>\overline{u}^{2}(\varphi_{1}).$ (38)
The kink is supersonic from the point of view both of the region before and after it.
IV.4 The quasi-solitons within the shocks
We studied previously the kinks and the solitons in the absence of damping kogan2 . In Section IV.2 we have shown that the kinks exist also in the presence of damping. So what about the solitons?
Looking at the Fig. 4 we understand that in the absence of damping ($R_{J}=\infty\rightarrow\gamma=0$) the motion of the fictitious particle starting and ending at $\varphi=\varphi_{1}$ is possible. This corresponds to the boundary conditions (23) with $\varphi_{2}=\varphi_{1}$ (the soliton). When we switch on the damping the solitons don’t exist any more, because the particle can’t return to its initial position. However, for weak damping, the particle after leaving the initial equilibrium position $\varphi_{1}$ will nearly return to this position after reflection from the opposite wall of the potential well (see Fig. 4). The part of the shock described by that motion will look very much like the soliton. The point of the reflection $\varphi_{0}$ can be found from the equation
$\Pi(\sin\varphi_{0})\approx\Pi(\sin\varphi_{1}).$ (39)
Substituting the formula for the potential energy (31) into (39) we can connect between $\varphi_{2}$ and $\varphi_{0}$, thus obtaining an alternative to (9) expression for the wave speed
$\overline{U}^{2}(\varphi_{1},\varphi_{0})=2\frac{\cos\varphi_{1}-\cos\varphi_{0}+(\varphi_{1}-\varphi_{0})\sin\varphi_{1}}{(\varphi_{1}-\varphi_{0})^{2}}.$ (40)
Note that the leading peak will be followed by the other ones, only (because of the continuous decrease of energy of the particle) the distance between the successive peaks will be decreasing and each successive peak will look less and less like the real soliton.
IV.5 Weak kinks and elementary weak shocks
For weak wave, characterized by the condition $\varphi_{1}-\varphi_{2}\ll 1$, the r.h.s. of (30) can be approximated as
$-\frac{d\Pi(\varphi)}{d\varphi}=\alpha(\varphi-\varphi_{1})(\varphi-\varphi_{2})(\varphi+\varphi_{3}),$ (41)
where
[…]
and (30) can be simplified to the damped Helmholtz-Duffing (dHD) equation
$\varphi_{\tilde{\tau}\tilde{\tau}}+\gamma\varphi_{\tilde{\tau}}=\alpha(\varphi-\varphi_{1})(\varphi-\varphi_{2})(\varphi+\varphi_{3}).$ (43)
For $\varphi_{2}<-\varphi_{3}$, Eq. (43) describes the kink, for $-\varphi_{3}<\varphi_{2}$ the shock.
Using the results of Appendix D we state that if the constants in (43) satisfy the condition
$\gamma=\sqrt{\frac{\alpha}{2}}(\varphi_{1}+\varphi_{2}+2\varphi_{3})=\sqrt{3\cos\overline{\varphi}}\tan\overline{\varphi},$ (44)
the solution of (43) satisfying the boundary conditions (23) is
$\varphi(\tilde{\tau})=\overline{\varphi}+\frac{\Delta\varphi}{2}\tanh\left(\beta\tilde{\tau}\right),$ (45)
where $\Delta\varphi=\varphi_{1}-\varphi_{2}$ and
$\beta=\sqrt{\frac{\alpha}{8}}\Delta\varphi=\sqrt{\frac{\cos\overline{\varphi}}{48}}\Delta\varphi.$ (46)
Note that for $\varphi_{2}<-\varphi_{3}$, Eq. (44) is both the condition of the kink existence (fine tuning we talked about previously) and of the kink being given by the elementary function. The shock doesn’t demand find tuning, so for $-\varphi_{3}<\varphi_{2}$, Eq. (44) is only the condition of the shock being given by the elementary function.
If the assumption $\Delta\varphi\ll 1$ is strengthened to $\Delta\varphi\ll\tan\overline{\varphi}$, Eq. (43) can be approximated as
$\varphi_{\tilde{\tau}\tilde{\tau}}+\gamma\varphi_{\tilde{\tau}}=\frac{\sin\overline{\varphi}}{2}(\varphi-\varphi_{1})(\varphi-\varphi_{2}).$ (47)
It is shown in Appendix D the solution of (47) satisfying the boundary conditions (23) is
$\varphi(\tilde{\tau})=\varphi_{2}+\frac{\Delta\varphi}{\left[\exp(\beta^{\prime}\tilde{\tau})+1\right]^{2}},$ (48)
where
$\beta^{\prime}=\sqrt{\Delta\varphi\sin\overline{\varphi}/12},\hskip 28.45274pt\gamma=5\beta^{\prime}.$ (49)
In the vicinity of $\varphi=0$ and for $\Delta\varphi\ll\tan\overline{\varphi}$, the condition (44) corresponds to strongly overdamped oscillator, and the condition (49), to slightly overdamped oscillator. The shock described by Eq. (47) exists for any value of $\gamma$. Equation (49) is only the condition of it being given by the elementary function.
In both particular cases studied above, because of the overdamped nature of the oscillations, the shock wave is monotonic. Note that the detailed study of the shock wave structure in a strongly nonlinear lattice with viscous dissipation was presented in Ref. nesterenko where, in particular, the dichotomy between the oscillatory and the monotonic shock waves was analysed quantitatively. We, however, postpone a similar analysis for the shocks in the JTL until the next publication.
It is interesting to see how the results obtained above can be recovered in the framework of the approach based on Eq. (36). If $\varphi_{1}$ and $\varphi_{2}$ are close enough to each other, we can approximate $\cos\varphi$ between $\varphi_{1}$ and $\varphi_{2}$ as a second degree polynomial in $\varphi$
$\cos\varphi=\cos\overline{\varphi}-\varphi^{\prime}\sin\overline{\varphi}-\frac{\varphi^{\prime 2}}{2}\cos\overline{\varphi},$ (50)
where $\varphi^{\prime}=\varphi-\overline{\varphi}$. The boundary conditions (23) obviously give
$E\left(\varphi_{1,2}\right)=0,$ (51)
thus we may try the solution of (36) in the form
$E(\varphi)=\psi\left(\varphi-\varphi_{1}\right)\left(\varphi-\varphi_{2}\right),$ (52)
where $\psi$ is a constant. Substituting (52) into (36) we obtain that (52) is indeed the solution, provided Eq. (44) is valid and $\psi^{2}=\cos\overline{\varphi}/12$. To find $\varphi(\tilde{\tau})$ we have to solve equation
$\frac{d\varphi}{d\tilde{\tau}}=\sqrt{\frac{\cos\overline{\varphi}}{12}}\left(\varphi-\varphi_{1}\right)\left(\varphi-\varphi_{2}\right).$ (53)
The solution is Eq. (45).
When $\Delta\varphi\ll\overline{\varphi}$ we can keep in the r.h.s. of (50) only the first two terms and present (36) (with the accepted precision) as
$\frac{1}{2}\frac{d^{2}E^{2}}{d\varphi^{2}}+\gamma\frac{dE}{d\varphi}=\overline{U}^{2}-\cos\overline{\varphi}+\left(\varphi-\varphi_{2}\right)\sin\overline{\varphi}.$ (54)
In this case the solution of (54) satisfying the boundary conditions Eq. (51) is
$E=\chi\left(\varphi-\varphi_{2}\right)\left[\left(\varphi-\varphi_{2}\right)^{1/2}-\left(\Delta\varphi\right)^{1/2}\right].$ (55)
Substituting (55) into (54) and equating coefficients before the same powers of $\left(\varphi-\varphi_{2}\right)^{1/2}$ in the l.h.s and in the r.h.s. of the equation, we obtain (49) and $\chi^{2}=\sin\overline{\varphi}/3$. The function $E(\varphi)$ being found, we can find the function $\varphi(\tilde{\tau})$ by solving equation
$\frac{d\varphi}{d\tilde{\tau}}=\sqrt{\frac{\sin\overline{\varphi}}{3}}\left(\varphi-\varphi_{2}\right)\left[\left(\varphi-\varphi_{2}\right)^{1/2}-\left(\Delta\varphi\right)^{1/2}\right].$ (56)
The solution is given by (48).
V Conclusions
We considered the interaction of the sound waves with the shock waves in the JTL. The formulas for the reflection and the transmission coefficients turned out to be very simple and appealing.
We established the relation between the shocks existing in the lossy JTL and the kinks which, as we now understand, exist both in the lossy and in the lossless JTL. We also have shown that the profile of the shock in the lossy JTL demonstrates oscillatory behavior with leading peaks resembling solitary waves if the losses are weak.
We found that the nonlinear equation describing the weak kinks and the weak shocks can be integrated (in particular cases) in terms of elementary functions.
Appendix A The JTL composed of superconducting grains
A physically appealing model of the JTL composed of superconducting grains is presented in Fig. 8 (for simplicity we ignored the shunting capacitor).
Figure 8: Josephson transmission line composed of superconducting grains
Here, we take as the dynamical variables the phases at the grains $\Phi_{n}$ and the potentials of the grains $V_{n}$. The circuit equations are
$\frac{\hbar}{2e}\frac{d\Phi_{n}}{dt}$ $=v_{n}$ (57a)
$C\frac{dv_{n}}{dt}$ $=I_{c}\sin\left(\Phi_{n-1}-\Phi_{n}\right)-I_{c}\sin\left(\Phi_{n}-\Phi_{n+1}\right)$
$+\frac{1}{R_{J}}\left(v_{n-1}-2v_{n}+v_{n+1}\right).$ (57b)
We realise that Eq. (II) follows from Eq. (A) if we substitute $\varphi_{n}=\Phi_{n-1}-\Phi_{n}$. Also, if we exclude $v_{n}$ from (57a), (57b) we obtain
$\frac{d^{2}\Phi_{n}}{d\tau^{2}}$ $=$ $\sin\left(\Phi_{n-1}-\Phi_{n}\right)-\sin\left(\Phi_{n}-\Phi_{n+1}\right)$ (58)
$+$ $\frac{Z_{J}}{R_{J}}\frac{d}{d\tau}\left(\Phi_{n-1}-2\Phi_{n}+\Phi_{n+1}\right),$
which is the particular case of the Fermi-Pasta-Ulam-Tsingou equation (with losses).
It is interesting to compare (58) with the equation from Ref. malomed , describing the chain of interacting particles with friction
$m\frac{d^{2}y_{n}}{d\tau^{2}}=$displaystyle-- $\frac{\partial}{\partial y_{n}}\left[U\left(y_{n-1}-y_{n}\right)+U\left(y_{n+1}-y_{n}\right)\right]$ (59)
-- $\gamma\frac{dy_{n}}{d\tau},$
where $y_{n}$ are displacements of particles in the chain and $U(z)$ is the potential of the interparticle interaction. Comparison shows different character of the losses in the systems.
It is also interesting to compare the one-dimensional Josephson-junction array, described by the discretized version of the perturbed sine-Gordon equation ustinov
$\frac{d^{2}\varphi_{n}}{d\tau^{2}}+\gamma\frac{d\varphi_{n}}{d\tau}+\sin\varphi_{n}$displaystyle-- $\frac{1}{a^{2}}\left(\varphi_{n-1}-2\varphi_{n}+\varphi_{n+1}\right)$ (60)
$=$ $I/I_{c},$
with the equation obtained by excluding $v_{n}$ from (II)
$\frac{d^{2}\varphi_{n}}{d\tau^{2}}=\sin\varphi_{n+1}-2\sin\varphi_{n}+\sin\varphi_{n+1}$
$+\left(\frac{Z_{J}}{R_{J}}+\frac{C_{J}}{C}\frac{d}{d\tau}\right)\frac{d}{d\tau}\left(\varphi_{n+1}-2\varphi_{n}+\varphi_{n+1}\right)$ (61)
shows that the nature of nonlinearity in the systems is different. Neither does (A) in the continuum approximation coincides with the sine-Gordon equation with losses landauer .
Appendix B The continuum and the discrete JTL
Natural question is how good is the continuum approximation used everywhere in this paper? To answer this question let us focus on Eq. (A). The continuum approximation consists in promoting the discrete variable $n$ to the continuous variable $Z$ and approximating the discrete second order derivatives in the r.h.s. of (A) by the continuous derivatives:
$\sin\varphi_{n+1}-2\sin\varphi_{n}+\sin\varphi_{n+1}$ $=\frac{\partial^{2}\sin\varphi}{\partial Z^{2}}$ (62a)
$\varphi_{n+1}-2\varphi_{n}+\varphi_{n+1}$ $=\frac{\partial^{2}\varphi}{\partial Z^{2}}.$ (62b)
To find the limits of applicability of this approximation, let us go one step further and consider the quasi-continuum approximation kogan2
$\sin\varphi_{n+1}-2\sin\varphi_{n}$ $+$ $\sin\varphi_{n+1}=\frac{\partial^{2}\sin\varphi}{\partial Z^{2}}+\frac{1}{12}\frac{\sin\partial^{4}\varphi}{\partial Z^{4}}.$
In this approximation, the equation describing the localized travelling wave is
$\frac{1}{12\overline{U}^{2}}\frac{d^{2}\sin\varphi}{d\tau^{2}}+\frac{C_{J}}{C}\frac{d^{2}\varphi}{d\tau^{2}}+\frac{Z_{J}}{R_{J}}\frac{d\varphi}{d\tau}=\overline{U}^{2}\varphi-\sin\varphi+F$
(compare with (27)). So the continuum approximation is applicable if either $C_{J}/C\gg 1$ or $Z_{J}/R_{J}\gg 1$.
Lossless JTL clearly corresponds to $R_{J}=\infty$. In addition, while talking about the lossless system in the present paper, we had in mind Eq. (B) containing only the second term in the l.h.s.. Previously kogan2 , we considered unshunted JTL in the quasi-continuum approximation, that is Eq. (B) with only the first term in the l.h.s.. However, because of the similarity of the terms, the kinks and the solitons obtained in the framework of these two considerations are qualitatively very similar. (The quantitative differences do exist. Thus the equation for the ”soliton” velocity (40) is different from the equation for the soliton velocity (36) from Ref. kogan2 ).
Appendix C Numerical integration of Eq. (27)
To integrate numerically Eq. (27) (for the sake of definiteness we consider the case $\varphi_{1}>0$), we should present it as an equation with the initial conditions. To do it let us first concentrate on the vicinity of the point $\varphi_{1}$. There the equation can be linearised and presented as
$\frac{d^{2}\varphi}{d\tilde{\tau}^{2}}+\gamma\frac{d\varphi}{d\tilde{\tau}}+\left(\cos\varphi_{1}-\overline{U}_{\varphi_{2},\varphi_{1}}^{2}\right)(\varphi-\varphi_{1})=0.$
Solving Eq. (C) while taking into account first of the boundary conditions (23) we obtain
$\varphi(\tilde{\tau})=\varphi_{1}-Ae^{\kappa\tilde{\tau}},$ (66)
where
$\kappa=\sqrt{\gamma^{2}/4-\cos\varphi_{1}+\overline{U}_{\varphi_{2},\varphi_{1}}^{2}}-\gamma/2$ (67)
and $A$ is an arbitrary positive constant.
Let us put in (66) $\tilde{\tau}=0$. We obtain
$\varphi(0)=\varphi_{1}-A$ (68)
(for (68) to make sense, the constant $A$ should satisfy the condition $A\ll\varphi_{2}$). Differentiating (66) with respect to $\tilde{\tau}$ and again putting $\tilde{\tau}=0$ we obtain
$\left.\frac{d\varphi}{d\tilde{\tau}}\right|_{\tau=0}=-A\kappa.$ (69)
Equations (68) and (69) we use as the initial conditions while numerically solving (27) for $\tau>0$.
Appendix D The elementary particular solutions of the generalized dHD equation
We consider the generalized dHD equation
$x_{\tau\tau}+\gamma x_{\tau}=\alpha x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right),$ (70)
where $n$ is an integer, $\gamma$, $\alpha$, $x_{1},x_{3}$ are positive constants.
Looking for an elementary particular solution of (70) we try first to integrate the equation by quadrature. To do it let us get read of the first derivative with respect to $\tau$ in (70) by introducing new dependent variable
$x(\tau)=e^{-m\tau}w(\tau),$ (71)
where the parameter $m$ will be determined later. This change of variable turns Eq. (70) into
$w_{\tau\tau}+(\gamma-2m)w_{\tau}=\alpha e^{-2nm\tau}w^{2n+1}$
$-\alpha(x_{1}-x_{3})e^{-nm\tau}w^{n+1}-\left(m^{2}-m\gamma+\alpha x_{1}x_{3}\right)w.$ (72)
We see that the choice $m=\gamma/2$ would cancel the first derivative term in the equation, but the price is too high - the coefficients before the nonlinear terms in the equation would become explicitly $\tau$-dependent.
However we can achieve our aim in a different way. Let us first kill the last term in the r.h.s. of (D) by choosing $m$ satisfying the equation
$m^{2}-m\gamma+\alpha x_{1}x_{3}=0.$ (73)
Let us also make the change of the independent variable
$\tau\to z(\tau)=e^{-nm\tau}.$ (74)
After that, Eq. (D) becomes
$n^{2}m^{2}w_{zz}+n\left[(n+1)m^{2}-\alpha x_{1}x_{3}\right]\frac{w_{z}}{z}$
$=\alpha w^{2n+1}-\alpha(x_{1}-x_{3})\frac{w^{n+1}}{z}.$ (75)
D.1 Integration by quadrature
Consider the particular case $x_{3}=x_{1}$. In this case (D) can be easily integrated by quadrature, provided the coefficient before the first derivative is equal to zero, i.e. $m$ is
$m=\sqrt{\frac{\alpha}{n+1}}x_{1}.$ (76)
Note that the condition of compatibility of (76) and (73) (the latter with $x_{3}=x_{1}$) is
$\gamma=(n+2)m.$ (77)
The condition being assumed, Eq.(D) becomes
$w_{zz}-\frac{n+1}{n^{2}x_{1}^{2}}w^{2n+1}=0$ (78)
(compare (78) with (70)). Multiplying Eq.(78) by $w_{z}$ and integrating with respect to $z$ we have
$w_{z}^{2}-\frac{1}{n^{2}x_{1}^{2}}w^{2n+2}=2E,$ (79)
where $E$ is the integration constant.
D.2 The elementary solutions
Let us demand that the solution of (70) satisfies the boundary conditions
$\lim_{\tau\to-\infty}x(\tau)=x_{1}^{1/n},\hskip 28.45274pt\lim_{\tau\to+\infty}x(\tau)=0.$ (80)
Note that if the solution of (70) $x(\tau)$ should exist and remain finite for $\tau\in(-\infty,+\infty)$, Eq. (80) are the only possible boundary conditions at $\tau=\pm\infty$ (apart from changing $x_{1}$ to $x_{3}$). The trajectory should start in the infinite past in the unstable fixed point and end in the infinite future in the stable fixed point.
From (80) follows that the solution of (79) should satisfy the boundary condition
$\lim_{z\to\infty}z^{1/n}w(z)=x_{1},$ (81)
and we should substitute $E=0$ into (79). This gives us the opportunity to integrate the equation not by quadrature, but in terms of elementary function:
$w(z)=\frac{x_{1}^{1/n}}{\left(1+z\right)^{1/n}}.$ (82)
Finally substituting (82) into (71), we obtain elementary particular solution of (70)
$x(\tau)=\frac{x_{1}^{1/n}.}{\left[\exp\left(nm\tau\right)+1\right]^{1/n}},$ (83)
which exists, provided the constants of (70) are connected by the relation (77).
Now we have a pleasant surprise: the elementary solution (82) solves (D) also when $x_{3}\neq x_{1}$. In fact, substituting (82) into (D) we obtain
$\frac{(n+1)m^{2}}{(1+z)^{2}}-\frac{(n+1)m^{2}-\alpha x_{1}x_{3}}{z(1+z)}$
$=\frac{\alpha x_{1}^{2}}{(1+z)^{2}}-\frac{\alpha(x_{1}-x_{3})x_{1}}{z(1+z)},$ (84)
and mysteriously the terms in the l.h.s and in the r.h.s. of the equation are equal pairwise, provided $m$ is given by (76). Hence (83) is the elementary solution of (70) valid also when $x_{3}\neq x_{1}$). The condition of compatibility of (73) and (76) in the general case is
$\gamma=\sqrt{\frac{\alpha}{n+1}}[x_{1}+(n+1)x_{3}]$ (85)
(compare with(77)).
D.3 The strongly asymmetric case
For $x_{3}\gg x_{1}$ there exists additional elementary solution of (70). In this case the equation can be approximated as
$x_{\tau\tau}+\gamma x_{\tau}=\alpha x_{3}x\left(x^{n}-x_{1}\right)$
$=\alpha x_{3}x\left(x^{n/2}-x_{1}^{1/2}\right)\left(x^{n/2}+x_{1}^{1/2}\right).$ (86)
Comparing (D.3) with (70) (and keeping in mind the elementary solution of the latter (83), we obtain the elementary solution of (D.3)
$x(\tau)=\frac{x_{1}^{1/n}}{\left[\exp\left(nm^{\prime}\tau\right)+1\right]^{2/n}}.$ (87)
where
[…]
D.4 Abel equation
The obtained elementary solutions become more transparent after we notice that introducing new dependent variable $p$
$p=x_{\tau}$ (89)
and considering $x$ as the independent variable, we may write down (70) as Abel equation of the second kind polyanin
$pp_{x}+\gamma p=\alpha x\left(x^{n}-x_{1}\right)\left(x^{n}+x_{3}\right).$ (90)
For $\gamma$ and $\alpha$ connected by the formula (85), Eq. (90) has the particular solution
$p=\sqrt{\frac{\alpha}{n+1}}x\left(x^{n}-x_{1}\right)$ (91)
Substituting $p(x)$ into (89) and integrating thus obtained differential equation we obtain (83).
Approximating (90) for $x_{3}\gg x_{1}$ as
$pp_{x}+\gamma p$ $=$ $\alpha x_{3}x\left(x^{n}-x_{1}\right)$
$=$ $\alpha x_{3}x\left(x^{n/2}-x_{1}^{1/2}\right)\left(x^{n/2}+x_{1}^{1/2}\right),$
we obtain from (91) the particular solution of (D.4)
$p=\sqrt{\frac{2\alpha x_{3}}{n+2}}x\left(x^{n/2}-x_{1}^{1/2}\right),$ (93)
and from (85) the connection between $\gamma$ and $\alpha$ in the present case; the latter turns out to be identical to that given by (88). Substituting obtained $p(x)$ into (89) and integrating thus obtained differential equation we obtain (87).
Acknowledgements.
I am grateful to A. Abanov, M. Goldstein, B. Malomed and J. Cuevas-Maraver for the discussion. I am also very grateful to Donostia International Physics Center (DIPC) for the hospitality during my visit, when the paper was finalised.